# Hashing-Baseline: Rethinking hashing in the age of pretrained models

### v1

@@ Abstract
Information retrieval with compact binary embeddings, also referred to as hashing, is crucial for scalable fast search applications, yet state-of-the-art hashing methods require expensive, scenario-specific training. In this work, we introduce Hashing-Baseline, a strong training-free hashing method leveraging powerful pre-trained encoders that produce rich pretrained embeddings. We revisit classical, training-free hashing techniques—principal component analysis, random orthogonal projection, and threshold binarization—to produce a strong baseline for hashing. Our approach combines these techniques with frozen embeddings from state-of-the-art vision and audio encoders to yield competitive retrieval performance without any additional learning or fine-tuning. To demonstrate the generality and effectiveness of this approach, we evaluate it on standard image retrieval benchmarks as well as a newly introduced benchmark for audio hashing. [^1] Code is released at: https://github.com/ilyassmoummad/hashing-baseline
Index Terms—  Hashing baseline, image retrieval, audio retrieval, binary codes, pre-trained encoders.
@@ 3.4 Theoretical Motivation
Hashing-Baseline is grounded in classical results: (i) PCA reduces information redundancy and noise in the pre-trained embeddings, and retains only the most informative dimesions, while preserving geometry in the reduced space, as ensured by the Johnson–Lindenstrauss lemma [9]; (ii) random orthogonal projections redistribute the concentrated variance across bits; (iii) random hyperplane hashing [10] links Hamming distance to angular similarity.

### v2

@@ Abstract
Information retrieval with compact binary codes, also referred to as hashing, is crucial for scalable fast search applications, yet state-of-the-art hashing methods require expensive, scenario-specific training. In this work, we introduce Hashing-Baseline, a strong training-free hashing method leveraging powerful pre-trained encoders that produce rich embeddings. We revisit classical, training-free hashing techniques—principal component analysis, random orthogonal projection, and threshold binarization—to produce a strong baseline for hashing. Our approach combines these techniques with frozen embeddings from state-of-the-art vision and audio encoders to yield competitive retrieval performance without any additional learning or fine-tuning. To demonstrate the generality and effectiveness of this approach, we evaluate it on standard image retrieval benchmarks as well as a newly introduced benchmark for audio hashing. [^1] Code is released at: https://github.com/ilyassmoummad/hashing-baseline
Index Terms—  Hashing baseline, image retrieval, audio retrieval, binary codes, pre-trained models.
@@ 3.4 Theoretical Motivation
Hashing-Baseline is grounded in classical results: (i) PCA reduces information redundancy and noise in the pre-trained embeddings, and retains only the most informative dimensions, after which random projections preserve geometry in the reduced space with high probability, as ensured by the Johnson–Lindenstrauss lemma [10]; (ii) random orthogonal projections redistribute the concentrated variance across bits; (iii) random hyperplane hashing [15] links Hamming distance to angular similarity.
@@ 5 Conclusion & Perspectives
6 Acknowledgment
This work was supported by the Pl@ntAgroEco project, funded by the ”Agence Nationale de la Recherche” (ANR) under the France 2030 program, within the “Agroécologie et Numérique” initiative (reference ANR-22-PEAE-0009). The authors gratefully acknowledge this support.